# Bootstrapped Control Limits for Score-Based Concept Drift Control Charts

### v1

@@ 3 Nested Bootstrap Procedure for Computing the CL
Suppose the training sample $\mathcal{D}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{n}$ is drawn i.i.d. from some joint distribution $\mathbb{P}_{0}(Y,\mathbf{X})$ for which the conditional distribution $\mathbb{P}(Y|\mathbf{X};\boldsymbol{\theta}^{0})$ can be implicitly represented by the parametric family $g_{\gamma}(\boldsymbol{\theta};\mathbf{x})$ of supervised learning models. Let $\mathbb{E}_{0}[\cdot]$ denote the expectation operator with respect to $\mathbb{P}_{0}(Y,\mathbf{X})$.
We apply the supervised learning model to the training dataset $\mathcal{D}$ and compute the score vectors. For notational simplicity, we denote these score vectors as $\mathcal{S}=\{\mathbf{s}_{i}:i=1,2,\cdots,n\}$ instead of using the functional notation $\mathbf{s}(\hat{\boldsymbol{\theta}};\mathbf{x}_{i},y_{i})$ from previous sections. Recall, each $\mathbf{s}_{i}$ is the derivative of the component of the model fitting objective function (the negative log-likelihood or a regularized version, with no optimization constraints) associated with observation $(\mathbf{x}_{i},y_{i})$. Let $\overline{\mathbf{s}}=1/n\sum_{i=1}^{n}\mathbf{s}_{i}$ and $\widehat{\boldsymbol{\Sigma}}=1/n\sum_{i=1}^{n}(\mathbf{s}_{i}-\overline{\mathbf{s}})(\mathbf{s}_{i}-\overline{\mathbf{s}})^{\top}$ denote the sample mean vector and covariance matrix of $\mathcal{S}$, respectively. By construction of the estimator $\hat{\boldsymbol{\theta}}$, $\overline{\mathbf{s}}=0$.
Now we consider a set of new observations $\mathcal{D}^{\text{new}}=\{(\mathbf{x}_{n+i},y_{n+i}):i=1,2,\cdots\}$ drawn i.i.d. from the same distribution as $\mathcal{D}$, to represent the situation that there is no shift in the predictive distribution. The score vector $\mathbf{s}_{n+i}=\mathbf{s}(\hat{\boldsymbol{\theta}};\mathbf{x}_{n+i},y_{n+i})$ for each new observation depends on $(\mathbf{x}_{n+i},y_{n+i},\hat{\boldsymbol{\theta}})$. Let $\mathcal{S}^{\text{new}}=\{\mathbf{s}_{n+i}:i=1,2,\cdots\}$ denote the new score vectors, and note that $\mathcal{S}^{\text{new}}$ depends on $\mathcal{D}$ only via the fitted model parameters $\hat{\boldsymbol{\theta}}$. Although the regularization parameters $\mathbf{\gamma}$ are also estimated from $\mathcal{D}$, we treat them as fixed for tractability and use the same values when fitting all models within the bootstrapping procedure described below. Consequently, conditioned on $\hat{\boldsymbol{\theta}}$, $\mathcal{S}^{\text{new}}$ constitutes an i.i.d. sample with each $\mathbf{s}_{n+i}$, $i=1,2,\cdots$, having some common mean $\mathbf{\mu}(\hat{\boldsymbol{\theta}})$ and covariance matrix $\mathbf{V}(\hat{\boldsymbol{\theta}})$ that are deterministic functions of $\hat{\boldsymbol{\theta}}$, which we denote by
@@
Although $\overline{\mathbf{s}}=\mathbf{0}$, it is not the case that $\mathbf{\mu}(\hat{\boldsymbol{\theta}})=\mathbf{0}$, because of estimation error in $\hat{\boldsymbol{\theta}}$ and because $\mathbf{s}_{n+i}$ are computed for new observations that are independent of the training data $\mathcal{D}$ to which $g_{\gamma}(\hat{\boldsymbol{\theta}};\mathbf{x})$ is fit. For similar reasons, $\widehat{\boldsymbol{\Sigma}}$ should not be viewed as an estimator of $\mathbf{V}(\hat{\boldsymbol{\theta}})$.
@@
$\mathbb{E}_{0}\!\left[\mathbf{z}_{n+i}\,\middle|\,\hat{\boldsymbol{\theta}}\right]$ $=\bigl{[}1-(1-\lambda)^{i}\bigr{]}\,\boldsymbol{\mu}(\hat{\boldsymbol{\theta}}),\qquad\operatorname{Cov}_{0}\!\left[\mathbf{z}_{n+i}\,\middle|\,\hat{\boldsymbol{\theta}}\right]$ $=\frac{\lambda}{2-\lambda}\,\bigl{[}1-(1-\lambda)^{2i}\bigr{]}\,\mathbf{V}(\hat{\boldsymbol{\theta}}).$ (3.2)
@@
Input: the full training sample $\mathcal{D}$; the desired false-alarm probability $\alpha$; and the MEWMA parameter $\lambda$.
@@
1) Fit and tune a supervised learning model $g_{\gamma}(\hat{\boldsymbol{\theta}};\mathbf{x})$ to $\mathcal{D}$ using CV to select hyperparameters $\gamma$.
@@
Within the context of the nested bootstrap procedure, each outer bootstrap sample $\mathcal{D}^{b}$ in Step 3-I assumes the role of $\mathcal{D}$, and will account for variability in $\hat{\boldsymbol{\theta}}$ that results from fitting the model to the training data, and its OOB sample $\mathcal{D}_{\text{OOB }}^{b}$ assumes the role of $\mathcal{D}^{\text{new }}$ and will account for variability in the future data. A single (non-nested) boostrap procedure would only account for the latter and underestimate the CL. Let $n^{b}$ denote the number of unique observations in $\mathcal{D}$ that were sampled in $\mathcal{D}^{b}$, in which case $\left|\mathcal{D}_{\text{OOB }}^{b}\right|=n-n^{b}$. By the 0.632 rule, unless $n$ is exceptionally small, we are guaranteed that $n^{b}\cong 0.632n$, in which case $|\mathcal{D}_{OOB}^{b}|\cong 0.368n$. For notational simplicity, in the following derivations we assume each $n^{b}=0.632n$.

### v2

@@ 3 Nested Bootstrap Procedure for Computing the CL
Suppose the training sample $\mathcal{D}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{n}$ is drawn i.i.d. from some joint distribution $\mathbb{P}_{0}(Y,\mathbf{X})$ for which the conditional distribution $\mathbb{P}(Y|\mathbf{X};\boldsymbol{\theta}^{0})$ can be implicitly represented by the parametric family $g_{\gamma}(\boldsymbol{\theta};\mathbf{x})$ of supervised learning models, where $\gamma$ denotes fixed hyperparameters or regularization constants. Let $\mathbb{E}_{0}[\cdot]$ denote the expectation operator with respect to $\mathbb{P}_{0}(Y,\mathbf{X})$.
We apply the supervised learning model to the training dataset $\mathcal{D}$ and compute the score vectors. For notational simplicity, we denote these score vectors as $\mathcal{S}=\{\mathbf{s}_{i}:i=1,2,\cdots,n\}$ instead of using the functional notation $\mathbf{s}(\hat{\boldsymbol{\theta}};\mathbf{x}_{i},y_{i})$ from previous sections. Recall, each $\mathbf{s}_{i}$ is the derivative of the component of the model fitting objective function (the negative log-likelihood or a regularized version, with no optimization constraints) associated with observation $(\mathbf{x}_{i},y_{i})$. Let $\overline{\mathbf{s}}=1/n\sum_{i=1}^{n}\mathbf{s}_{i}$ and $\widehat{\boldsymbol{\Sigma}}=1/n\sum_{i=1}^{n}(\mathbf{s}_{i}-\overline{\mathbf{s}})(\mathbf{s}_{i}-\overline{\mathbf{s}})^{\top}$ denote the sample mean vector and covariance matrix of $\mathcal{S}$, respectively. By construction, $\overline{\mathbf{s}}=0$.
[…]
@@
[…]
@@
[…]
@@
It is worth noting that standard control charting often targets a specific in-control Average Run Length (ARL). However, in the context of score-based monitoring with complex models, the monitoring statistic is not generated from a stationary process with known parameters, but depends on the finite-sample variability of the fitted model. Furthermore, calculating ARLs via Monte Carlo simulation for our approach is computationally prohibitive due to the nested bootstrap required at each time step. Therefore, Algorithm 1 is designed to control the pointwise false-alarm rate $\alpha$ at each monitoring instance $i$. This ensures that the probability of a false alarm at any specific time $t$ is effectively bounded by $\alpha$, a property validated empirically in Section 4.
Input: the full training sample $\mathcal{D}$; user’s desired false-alarm rate $\alpha$; and the MEWMA parameter $\lambda$.
@@
1) Fit and tune a supervised learning model $g_{\gamma}(\hat{\boldsymbol{\theta}};\mathbf{x})$ to $\mathcal{D}$ using cross validation (CV) to select hyperparameters $\gamma$.
@@
Within the nested bootstrap procedure, each outer bootstrap sample $\mathcal{D}^{b}$ in Step 3-I assumes the role of the original training sample $\mathcal{D}$ and accounts for Phase I (model-estimation) uncertainty arising from fitting the supervised learning model, i.e., variability in $\hat{\boldsymbol{\theta}}$. The corresponding out-of-bag (OOB) sample $\mathcal{D}_{\mathrm{OOB}}^{b}$ plays the role of the future data stream $\mathcal{D}^{\mathrm{new}}$ and captures Phase II (prospective monitoring) variability in incoming observations. As a result, Algorithm 1 is designed to control the marginal pointwise false-alarm rate at each monitoring time, explicitly accounting for both Phase I uncertainty (via the outer bootstrap loop) and Phase II data variability (via the inner bootstrap loop). In contrast, a single (non-nested) bootstrap procedure would account only for Phase II variability and therefore substantially underestimate the control limit. Let $n^{b}$ denote the number of unique observations in $\mathcal{D}$ sampled in $\mathcal{D}^{b}$, in which case $\lvert\mathcal{D}_{\mathrm{OOB}}^{b}\rvert=n-n^{b}$.
We employ standard bootstrap sampling with replacement. Consequently, each bootstrap sample $\mathcal{D}^{b}$ contains approximately $1-(1-1/n)^{n}\approx 0.632$ fraction of unique observations from $\mathcal{D}$ (Efron 1983). The remaining $\approx 0.368n$ observations form the out-of-bag (OOB) sample $\mathcal{D}_{OOB}^{b}$, which serves as a proxy for independent future data. For notational simplicity, in the following derivations we assume each $n^{b}=0.632n$.